# Uncertainties from metrology in the integrated luminosity measurement with the updated design of a detector at CEPC

### v2

@@ 1 Introduction
In order to achieve precision goals of the electroweak physics program at CEPC at the $Z^{0}$ pole, integrated luminosity $\mathcal{L}$ should be known with the total relative uncertainty of order of $10^{-4}$ CEPC_CDR . In the CEPC Conceptual Design Report CEPC_CDR , beam properties and detector concept are discussed, with particular emphasis on the machine-detector interface (MDI) and integrated luminosity measurement. The current design of MDI at CEPC has a distinctive feature with respect to the proposal in CEPC_CDR - the luminometer is proposed to be placed at the z-axis. Although the impact of metrology on the integrated luminosity measurement at CEPC with the luminometer conventionally placed at the outgoing beams (s-axis) has been discussed in JINST , it requires revisiting for the newly proposed geometry, since the integrated luminosity precision critically depends on the luminometer’s position, as well as on the event selection.
@@
The paper is organized as follows: in Section 2 we discuss the current layout of the very forward region of CEPC. In Section 3 we review the impact of detector manufacturing, positioning and alignment (Section 3.1), as well as the impact of uncertainties of the knowledge of the beam properties (asymmetry of beam energies, beam energy spread - BES and beam synchronization) in section 3.2. Concluding discussion is given in Section 4.
@@ 2 Forward region at CEPC
The MDI region of CEPC is proposed to longitudinally cover the area of 12 m with the centrally positioned interaction point (IP). The components of the accelerator (without shielding) will be placed in an 118 mrad conus inside the detector, as illustrated in Fig. 1 JINST . The crossing angle in the horizontal plane between the colliding beams at the IP is 33 mrad. The final focus length of colliding beams is 220 cm CEPC_CDR . It is proposed that the luminometer at CEPC should be placed at 95 cm distance from the interaction point, covering the polar angle region from 30 mrad to 105 mrad, which corresponds to the luminometer aperture of 2.85 cm for the inner and 10.0 cm for the outer radius. The fiducial volume of luminometer should be between 53 mrad and 79 mrad. To prevent the leakage of the electromagnetic showers outside of the luminometer’s edges, an iron shielding of 5 mm thickness can be placed around the luminometer. However, this has been done under assumption that the luminometer will be placed around the outgoing beams and it is yet to be proven for the luminometer placed around the z-axis, since the symmetry of signal hits with respect to the detector axis will be lost, as illustrated in Fig. 2, where the hitmap in the first plane of the luminometer is presented.
The choice of technology of the luminometer at CEPC has not been finally determined yet. However, since this study is based on simulation of counting of low-angle Bhabha scattering without any other assumption but the angular acceptance of the luminometer, it is of no relevance for this study. If one would perform a more realistic simulation including detector response to signal and background, at least an energy cut on electron and positron tracks should be required to separate Bhabha events from the background. In such a case, effects like bias and resolution of energy measurement as well as the uncertainties from calibration of the luminometer should be included as additional sources of systematic uncertainty. The same happens if one includes other selection criteria based on polar and azimuthal angles of a Bhabha candidate.
@@ 3 Integrated luminosity measurement and systematic uncertainties
Further, feasibility and requirements for $\mathrm{10^{-4}}$ precision of counting at the $Z^{0}$ pole will be discussed, considering the uncertainty of each individual systematic effect contributes to the integrated luminosity at a level of $\mathrm{10^{-4}}$. In JINST , at one side of the luminometer the Bhabhas are counted in the full fiducial volume of the luminometer, while at the other side the inner radial acceptance of the fiducial volume is shrunken by 1 mm. In this study, the outer radial acceptance is also shrunken by 1 mm, as it has been done at OPAL OPAL , randomly to the left (L) and right (R) sides of the luminometer, on event by event basis. Described way of counting is called asymmetrical throughout the paper. For symmetric event selection the Bhabhas are counted in the full fiducial volumes of the both sides of luminometer. As will be discussed further, different detector positioning (s-axis and z-axis) interfere with a way of counting in a nontrivial way.
@@ 3.1 Uncertainties from mechanics and positioning
Figures 4-8 illustrate effects 1-4 respectively, for different ways of counting and different positions of the luminometer placed on the s- and z-axis.
As can be seen from Figures 4(b)-8(b), for the detector placed at the z-axis asymmetric counting performs similarly as counting in the full fiducial volume and also similarly as asymmetric counting for the detector placed at the s-axis. With the detector placed at the s-axis the difference between the two ways of counting is quite significant, as illustrated in Figures 4(a)-8(a).
@@
Figure 4: (a) Variations of the inner aperture, $\Delta r_{in}$, for two different ways of counting: in the full fiducial volume (S) and the LEP-style (AS), for detector placed on the s- and z-axis; (b) Count uncertainty as a function of $\Delta r_{in}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
@@
Figure 5: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for the RMS of the Gaussian spread of measured radial shower position with respect to the true one, $\sigma_{r}$. (b) Count uncertainty as a function of $\sigma_{r}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
@@
Figure 6: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal absolute uncertainty, $\Delta l$, of the longitudinal distance between left and right luminometer sides; (b) Count uncertainty as a function of $\Delta l$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
@@
Figure 7: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis against the RMS of a Gaussian distribution of radial fluctuations of the luminometer position with respect to the IP, $\sigma_{x_{IP}}$; (b) Count uncertainty as a function of $\sigma_{x_{IP}}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
@@
Figure 8: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis against the RMS of the Gaussian distribution of axial fluctuations of the luminometer position with respect to the IP, $\sigma_{z_{IP}}$; (b) Count uncertainty as a function of $\sigma_{z_{IP}}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
@@ 3.2 Beam related uncertainties
In the same manner as in the previous chapter, results for effects 1-3 are illustrated on Figures 9-12 and the upper limits, together with the upper limits of mechanical parameters from Section 3.1, are summarized in Table 2.
@@
As can be seen from Table 2, mechanical precision of the luminometer’s inner aperture mechanical precision seems to be the major challenge, since the Bhabha cross-section is scaling with the polar angle as $\sigma\sim 1/\theta^{3}$. The longitudinal boost of the Bhabha center-of-mass frame with respect to the laboratory frame, $\beta_{z}$, may be caused by any difference in beam energies, $\delta E$ ($\beta_{z}=2\cdot\delta E/\sqrt{s}$), in the form of bias, $\Delta E$, of one beam with respect to the other, or as a random asymmetry caused by the beam energy spread in the presence of other radiative losses (ISR and FSR). Limits of $\Delta E$ and $\sigma_{E_{BS}}$ given in Table 2 are derived from Figures 9(b) and 10(b) and rounded with respect to statistical size of the sample. From Table 2 it can be seen that in terms of the beam-related parameters precision, displacement of the luminometer from the s-axis to the z-axis will not impact the luminosity precision significantly, while some dependencies between observables will be modified due to loss of spacial symmetries of Bhabha events.
Figure 9: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal bias of energy of one beam with respect to the other, $\Delta E$; (b) Loss of the Bhabha count in the luminometer due to the longitudinal boost of the center-of-mass frame, $\beta_{z}(\Delta E)$, for two different ways of counting (S-triangles, AS-diamonds), with the luminometer on the z-axis.
Figure 10: (b) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal RMS of the BES Gaussian distribution, corresponding to the relative systematic uncertainty of $\mathcal{L}$ of $10^{-4}$, $\sigma_{E_{BS}}$; (b) Loss of the Bhabha count in the luminometer due to longitudinal boost of the Bhabha center-of-mass frame caused by BES, $\beta_{z}(\sigma_{E_{BS}})$, for two different ways of counting (S-triangles, AS-diamonds), with the luminometer on the z-axis.
Figure 11: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal radial IP position displacement with respect to the luminometer ($\Delta x_{IP}^{BS}$); (b) Count uncertainty as a function of $\Delta x_{IP}^{BS}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
Figure 12: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal axial IP position displacements with respect to the luminometer ($\Delta z_{IP}^{BS}$); (b) Count uncertainty as a function of $\Delta z_{IP}^{BS}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
@@ 4 Conclusion
Although the method of integrated luminosity has been studied in a great detail at LEP (LEP1 , LEP2 ), for each newly proposed $e^{+}e^{-}$ collider it is necessary to quantify the achievable luminosity precision, in particular at the $Z^{0}$ resonance. Here it is done for CEPC, taking into consideration the mechanical and beam-related requirements at the $Z^{0}$ pole, with the luminometer placed on the z-axis. It is the first attempt at a future Higgs factory to quantify systematic effects rising from metrology, with luminometer displaced from the the outgoing beams where it is conventionally positioned at colliders with a crossing angle.
@@
Although the simulated Bhabha samples are relatively small to be exact on boundaries of the considered systematics, distribution of central values presented in this study confirms that no effect seems to be more critical for precision measurement of the integrated luminosity if the luminometer is placed around the z-axis than if it is placed around the outgoing beams. The LEP-style asymmetric counting, however, will not be effective for the luminometer placed on the z-axis to cope with the effects arising from the left-right symmetry.
Complex metrology, together with the beam-related interactions affecting the Bhabha final state at small polar angles Voutsinas may call for consideration of an alternative or a complementary central process, like di-muon or di-photon production for the integrated luminosity measurement, with hopefully less complex systematic effects subjected to further studies ECFA . In addition, studies presented here should be complemented with a full simulation study of systematic effects arising from the detector design, technology and performance in the presence of realistic backgrounds.

### v3

@@ 1 Introduction
[…]
@@
The paper is organized as follows: in Section 2 we discuss the current layout of the very forward region of CEPC. In Section 3 we review the impact of detector manufacturing, positioning and alignment (Section 3.1), as well as the impact of uncertainties of the knowledge of the beam properties (asymmetry of beam energies, beam energy spread - BES and beam synchronization) in Section 3.2. The concluding discussion is given in Section 4.
@@ 2 Forward region at CEPC
The MDI region of CEPC is proposed to longitudinally cover the area of 12 m with the centrally positioned interaction point (IP). The components of the accelerator (without shielding) will be placed in an 118 mrad cone inside the detector, as illustrated in Fig. 1 Hou . The crossing angle in the horizontal plane between the colliding beams at the IP is 33 mrad. The final focus length of colliding beams is 220 cm CEPC_CDR . It is proposed that the luminometer at CEPC should be placed at 95 cm distance from the interaction point, covering the polar angle region from 30 mrad to 105 mrad, which corresponds to the luminometer aperture of 2.85 cm for the inner and 10.0 cm for the outer radius. The fiducial volume of luminometer should be between 53 mrad and 79 mrad. To prevent the leakage of the electromagnetic showers outside of the luminometer’s edges, an iron shielding of 5 mm thickness can be placed around the luminometer. However, this has been done under assumption that the luminometer will be placed around the outgoing beams and it is yet to be proven for the luminometer placed around the z-axis, since the symmetry of signal hits with respect to the detector axis will be lost, as illustrated in Fig. 2 where the hitmap in the first plane of the luminometer is presented.
@@
The choice of technology of the luminometer at CEPC has not been finally determined yet. However, since this study is based on simulation of counting of low-angle Bhabha scattering without any other assumption but the angular acceptance of the luminometer, it is of no relevance for this study. If one would perform a more realistic simulation including detector response to signal and background, at least an energy cut on electron and positron tracks should be required to separate Bhabha events from the background. In such a case, effects like bias and resolution of energy measurement as well as the uncertainties from calibration of the luminometer should be included as additional sources of systematic uncertainty. The same holds if one includes other selection criteria based on polar and azimuthal angles of a Bhabha candidate.
@@ 3 Integrated luminosity measurement and systematic uncertainties
Further, feasibility and requirements for $\mathrm{10^{-4}}$ precision of counting at the $Z^{0}$ pole will be discussed, considering that the uncertainty of each individual systematic effect contributes to the integrated luminosity at a level of $\mathrm{10^{-4}}$. In JINST , at one side of the luminometer the Bhabhas are counted in the full fiducial volume of the luminometer, while at the other side the inner radial acceptance of the fiducial volume is shrunken by 1 mm. In this study, the outer radial acceptance is also shrunken by 1 mm, as it has been done at OPAL OPAL , randomly to the left (L) and right (R) sides of the luminometer, on event by event basis. Described way of counting is called asymmetrical throughout the paper. For symmetric event selection the Bhabhas are counted in the full fiducial volumes of the both sides of luminometer. As will be discussed further, different detector positioning (s-axis and z-axis) interfere with a way of counting in a nontrivial way.
@@ 3.1 Uncertainties from mechanics and positioning
Figures 4-8 illustrate effects 1-4 respectively, for different ways of counting and different positions of the luminometer placed on the s- and z-axis. Values in histograms shown on Figures 4(a)-8(a) are based on the data from Figures 4(b)-8(b), obtained from the local fit in the region of interest. If the distribution on (b) was not symmetrical for positive and negative variations of the parameter, the more restrictive limit was considered.
Figures 4-8 show that for the detector placed at the z-axis asymmetric counting performs similarly as counting in the full fiducial volume, whereas with the detector placed at the s-axis the difference between the two ways of counting is quite significant. On Figure 4, which illustrates the impact of $\Delta r_{in}$, it can be seen that the counting uncertainties with the detector placed at the z-axis are similar to uncertainties when counting in the full fiducial volume with the detector placed at the s-axis. For $\sigma_{r}$ and $\sigma_{z_{IP}}$, the requirements are slightly more relaxed for the luminometer placed at the z-axis than with the asymmetric counting with the luminometer placed at the s-axis, as shown in Figures 5 and 8, whether for $\sigma_{x_{IP}}$ it’s the other way around. For $\Delta l$, the uncertainties are the same for both positions of the luminometer and for both ways of counting, as illustrated in Figure 6. However, one has to keep in mind that the results shown in this paper are based on distributions of mean values, each obtained from the simulated signal sample of $2\cdot 10^{7}$ events, with intrinsic relative statistical uncertainty of $\sim 2\cdot 10^{-4}$. Therefore, the extracted precision limits rather reflect the order of magnitude required to control certain mechanical or beam properties at the level of $10^{-4}$ than the exact values of detector or beam parameters.
@@
Figure 4: (a) Variations of the inner aperture (counting volume), $\Delta r_{in}$, for two different ways of counting: in the full fiducial volume (S) and the LEP-style (AS), for detector placed on the s- and z-axis; (b) Count uncertainty as a function of $\Delta r_{in}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@
Figure 5: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for the RMS of the Gaussian spread of measured radial shower position with respect to the true one, $\sigma_{r}$. (b) Count uncertainty as a function of $\sigma_{r}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@
Figure 6: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal absolute uncertainty, $\Delta l$, of the longitudinal distance between left and right luminometer sides; (b) Count uncertainty as a function of $\Delta l$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@
Figure 7: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis against the RMS of a Gaussian distribution of radial fluctuations of the luminometer position with respect to the IP, $\sigma_{x_{IP}}$; (b) Count uncertainty as a function of $\sigma_{x_{IP}}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@
Figure 8: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis against the RMS of the Gaussian distribution of axial fluctuations of the luminometer position with respect to the IP, $\sigma_{z_{IP}}$; (b) Count uncertainty as a function of $\sigma_{z_{IP}}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@ 3.2 Beam related uncertainties
In the same manner as in the previous chapter, results for effects 1-4 are illustrated on Figures 9-12 and the upper limits, together with the upper limits of mechanical parameters from Section 3.1, are summarized in Table 2. Each effect is considered to individually contribute to the the relative uncertainty of the integrated luminosity as $1\cdot 10^{-4}$.
Figure 9: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal bias of energy of one beam with respect to the other, $\Delta E$; (b) Loss of the Bhabha count in the luminometer due to the longitudinal boost of the center-of-mass frame, $\beta_{z}(\Delta E)$, for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
Figure 10: (b) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal RMS of the BES Gaussian distribution, corresponding to the relative systematic uncertainty of $\mathcal{L}$ of $10^{-4}$, $\sigma_{E_{BS}}$; (b) Loss of the Bhabha count in the luminometer due to longitudinal boost of the Bhabha center-of-mass frame caused by BES, $\beta_{z}(\sigma_{E_{BS}})$, for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
Figure 11: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal radial IP position displacement with respect to the luminometer ($\Delta x_{IP}^{BS}$); (b) Count uncertainty as a function of $\Delta x_{IP}^{BS}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
Figure 12: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal axial IP position displacements with respect to the luminometer ($\Delta z_{IP}^{BS}$); (b) Count uncertainty as a function of $\Delta z_{IP}^{BS}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@
As can be seen from Table 2, mechanical precision of the luminometer’s inner aperture seems to be the major challenge, since the Bhabha cross-section is scaling with the polar angle as $\sigma\sim 1/\theta^{3}$. The longitudinal boost of the Bhabha center-of-mass frame with respect to the laboratory frame, $\beta_{z}$, may be caused by any difference in beam energies, $\delta E$ ($\beta_{z}=2\cdot\delta E/\sqrt{s}$), in the form of bias of one beam with respect to the other, $\Delta E$, or as a random asymmetry caused by the beam energy spread or by other radiative losses. All limits of parameters given in Table 2 are derived from distributions of mean values in Figures 4(b) - 12(b), without considering statistical uncertainties from the limited size of the simulated samples. From Table 2 it can be seen that in terms of the beam-related parameters precision, displacement of the luminometer from the s-axis to the z-axis will not impact the luminosity precision significantly, while some dependencies between observables, especially $\Delta x_{IP}^{BS}$, $\Delta z_{IP}^{SY}$ (Figures 11 and 12) will be modified due to loss of spacial symmetries of Bhabha events.
@@ 4 Conclusion
Although the method of integrated luminosity has been studied in a great detail at LEP (LEP1 , LEP2 ), for each newly proposed $e^{+}e^{-}$ collider it is necessary to quantify the achievable luminosity precision, in particular at the $Z^{0}$ resonance. Here it is done for CEPC, taking into consideration the mechanical and beam-related requirements at the $Z^{0}$ pole, with the luminometer placed on the z-axis. It is the first attempt to quantify systematic effects rising from metrology with the luminometer displaced from the the outgoing beams, where it is conventionally positioned at colliders with a crossing angle.
@@
Although the simulated Bhabha samples are relatively small to be exact on boundaries of the considered systematics, distribution of central values presented in this study confirms that no effect seems to be more critical for precision measurement of the integrated luminosity if the luminometer is placed around the z-axis than if it is placed around the outgoing beams. The LEP-style asymmetric counting, however, will not be effective for the luminometer placed on the z-axis to cope with the effects arising from the left-right asymmetry.
Complex metrology, together with the beam-related interactions affecting the Bhabha final state at small polar angles Voutsinas may call for consideration of an alternative or a complementary central process for the integrated luminosity measurement, like di-photon or di-muon production (under the assumption that this process at the $Z^{0}$ pole energy doesn’t receive relevant contributions from the processes beyond the Standard Model), with hopefully less complex systematic effects subjected to further studies ECFA . In addition, studies presented here should be complemented with a full simulation study of systematic effects arising from the detector design, technology and performance in the presence of realistic backgrounds.